The input offset on the MC length servo board changes the lock point of the length loop (by how much? need to calibrate this slider into meters & Hz).
The SUM signal on the MC WFS is ~few 1000. This is several times larger than the pit/yaw signals. This is bad. it means that the TEM00 mode on the WFS (or what the WFS interperets as a TEM00) is larger than the TEM01/10 that its supposed to measure.
So if the beam moves on the WFS head it will convert this large common mode signal into a differential one.
We moved the MC Servo offset around from -3 to +3 V today and saw that it does affect the transmitted light level, but we need to think more to see how to put the offset at the real center of the resonance. This is complicated by the fact that the MCWFS loops seem to have some several minutes time constant so things are essentially always drifting.
I changed the McREFL SMOO to make it easier to use this noisy channel to diagnose small alignment changes:
caput C1:IOO-MC_RFPD_DCMON.SMOO 0.1
Huh? So should we ask them to put the container back? Or do you have some other theory about ETMX tripping that is not garbage related?
ETMX sus damping recovered.
Note: The giant metal garbage container was moved from the south west corner of CES months ago.
ETMX sus damping recovered. PSL enclousure is dusty at 20V rotation speed. Rainy days as outside condition.
It turned out the 'ringing' was caused by the respective other ETM still being aligned. For these reflection measurements both test masses of the other arm need to be misaligned. For the ETM it's sufficient to use the Misalign button in the medm screens, while the ITM has to be manually misaligned to move the reflected beam off the PD.
I did another round of armloss measurements today. I encountered some problems along the way
Each pair of readings gives the reflected power at the AS port normalized to the IMC stored power:
which is then averaged. The loss is calculated from the ratio of reflected power in the locked (L) vs misaligned (M) state from
Acquiring data this way yielded P_L/P_M=1.00507 +/- 0.00087 for the X arm and P_L/P_M=1.00753 +/- 0.00095 for the Y arm. With and (from m1=0.179, m2=0.226 and 91.2% and 86.7% mode matching in X and Y arm, respectively) this yields round trip losses of:
and , which is assuming a generalized 1% error in test mass transmissivities and modulation indices. As we discussed, this seems a little too good to be true, but at least the numbers are not negative.
Valve configuration: vacuum normal
Vacuum envelope: 23.5 C
RGA head: 46.6 C
c1psl finally booted up again, PMC and IMC are locked.
Trying to identify the hickup from the source code was fruitless. However, since the PMCTRANSPD channel acqusition failure occured long before the actual slow machine crashed, and since the hickup in the boot seemed to indicate a problem with daughter module identification, we started removing the DIO and DAQ modules:
The particle counter channel should be working again.
/cvs/cds/caltech/target/c1psl/psl.db lists the following channels for VMIVME3123:
Channels currently in use (and therefore not available in the medm screens):
Channels not currently in use (?):
There are plenty of channels available on the asynchronous ADC, so we could wire the relevant ones there if we done care about the 16 bit synchronous sampling (required for proper functionality?)
Alternatively, we could prioritize the Acromag upgrade on c1psl (DAQ would still be asynchronous, though). The PCBs are coming in next Monday and the front panels on Tuesday.
Some more info that might come in handy to someone someday:
The (nameless?) Windows 7 laptop that lives near MC2 and is used for the USB microscope was used for interfacing with c1psl. No special drivers were necessary to use the USB to RS232 adapter, and the RJ45 end of the grey homemade DB9 to RJ45 cable was plugged into the top port which is labeled "console 1". I downloaded the program "CoolTerm" from http://freeware.the-meiers.org/#CoolTerm, which is a serial protocol emulator, and it worked out of the box with the adapter. The standard settings fine worked for communicating with c1psl, only a small modification was necessary: in Options>Terminal make sure that "Enter Key Emulation" is set from "CR+LF" to "CR", otherwise each time 'Enter' is pressed it is actually sent twice.
Yes, that was one of the things that I wanted to look into. One thing Gautam and I did that I didn't mention was to reconnect the SRM satellite box and move the optic around a bit, which didn't change anything. Once the c1psl problem is fixed we'll resume with that.
The fringes seen on the oscope are mostly likely due to the interference from multiple light beams. If there are laser beams hitting mirrors which are moving, the resultant interference signal could be modulated at several Hertz, if, for example, one of the mirrors had its local damping disabled.
Speaking of which:
Using one of the grey RJ45 to D-Sub cables with an RS232 to USB adapter I was able to capture the startup log of c1psl (using the usb camera windows laptop). I also logged the startup of the "healthy" c1aux, both are attached. c1psl stalls at a point were c1aux starts testing for present vme modules and doesn't continue, however is not strictly hung up, as it still registers to the logger when external login attempts via telnet occur. The telnet client simply reports that the "shell is locked" and exits. It is possible that one of the daughter cards causes this. This seems to happen after iocInit is called by the startup script at /cvs/cds/caltech/target/c1psl/startup.cmd, as it never gets to the next item "coreRelease()". Gautam and I were trying to find out what happends inside iocInit, but it's not clear to us at this point from where it is even called. iocInit.c and compiled binaries exist in several places on the shared drive. However, all belong to R3.14.x epics releases, while the logfile states that the R3.12.2 epics core is used when iocInit is called.
Next we'll interrupt the autoboot procedure and try to work with the machine directly.
Using the PDA520 detector on the AS port I tried to get some better estimates for the round-trip loss in both arms. While setting up the measurement I noticed some strange output on the scope I'm using to measure the amount of reflected light.
The interferometer was aligned using the dither scripts for both arms. Then, ITMY was majorly misaligned in pitch AND yaw such that the PD reading did not change anymore. Thus, only light reflected from the XARM was incident of the AS PD. The scope was showing strange oscillations (Channel 2 is the AS PD signal):
For the measurement we compare the DC level of the reflection with the ETM aligned (and the arm locked) vs a misaligned ETM (only ITM reflection). This ringing could be observed in both states, and was qualitatively reproducible with the other arm. It did not show up in the MC or ARM transmission. I found that changing the pitch of the 'active' ITM (=of the arm under investigation) either way by just a couple of ticks made it go away and settle roughly at the lower bound of the oscillation:
In this configuration the PD output follows the mode cleaner transmission (Channel 3 in the screen caps) quite well, but we can't take the differential measurement like this, because it is impossible to align and lock the arm but them misalign the ITM. Moving the respective other ITM for potential secondary beams did not seem to have an obvious effect, although I do suspect a ghost/secondary beam to be the culprit for this. I moved the PDA520 on the optical table but didn't see a change in the ringing amplitude. I do need to check the PD reflection though.
Obviously it will be hard to determine the arm loss this way, but for now I used the averaging function of the scope to get rid of the ringing. What this gave me was:
(16 +/- 9) ppm losses in the x-arm and (-18+/-8) ppm losses in the y-arm
The negative loss obviously makes little sense, and even the x-arm number seems a little too low to be true. I strongly suspect the ringing is responsible and wanted to investigate this further today, but a problem with c1psl came up that shut down all work on this until it is fixed:
I found the PMC unlocked this morning and c1psl (amongst other slow machines) was unresponsive, so I power-cycled them. All except c1psl came back to normal operation. The PMC transmission, as recorded by c1psl, shows that it has been down for several days:
Repeated attempts to reset and/or power-cycle it by Gautam and myself could not bring it back. The fail indicator LED of a single daughter card (the DOUT XVME-212) turns off after reboot, all others stay lit. The sysfail LED on the crate is also on, but according to elog 10015 this is 'normal'. I'm following up that post's elog tree to monitor the startup of c1psl through its system console via a serial connection to find out what is wrong.
The microscope shipped back to the vendor for credit yesterday.
http://www.amscope.com/3-5x-180x-boom-stand-trinocular-zoom-stereo-microscope-with-144-led-ring-light-and-10mp-camera.html will be ordered today.
The actual unit we are getting has lockable zoom for better repeatability after calibration: SM-3NTPZZ-144
I've now made a DCC page for the mirror specifications, all revisions should be reflected there.
Over the last couple of days, I've been playing around with Rana's coating optimization code to come up with a coating design that will work for us. The basic idea is a to use MATLAB's particle swarm constrained optimization tool to minimize an error function that is a composite of four penalties:
On the AR side, I only considered 2 and 3. The weighting of these four components were set somewhat arbitrarily, but I seem to be able to get reasonable results so I am going with this for now.
From my first pass at it, the numbers I've been able to get, for 19 layer pairs, are (along with some plots):
(in this picture, the substrate is to the right of layer 38)
(substrate to the right of layer 38)
These numbers are already matching the specs we have on the DCC page currently. I am not sure how much better we can get the specs on the HR side keeping with 19 layer pairs...
All of this data, plus the code used to generate them, is on the gitlab coatings page...
Kyle took high quality images of the three sapphire prisms using the microscope @Downs. He analyzed the images to see the radius of the groove.
They all look sufficiently sharp for a 46um steel wire. Thumbs up.
I am curious to see how the wire Q is with grooved sapphires, ungrooved sapphires, grooved ruby, grooved aluminum stand off, and so on.
OK, but the questions still stands: "Assuming we want a 1% calibration at 50-500 Hz, what is the requirement on the frequency noise PSD curve?"
We get SNR in two ways: the amplitude of applied force and the integration time. So we are limited in two ways: stability of the lock to applied forces and time of locklosses / calibration fluctuations.
The MET#1 particle counter was moved from CES wall at ITMX to PSL enclousure south west corner at 11am.
The HEPA filter speed at the Variac was turned down to 20V from 40
This counter pumps air for 1 minute in every 20 minutes. Soft foam in bags used to minimize this shaking as it is clamped.
This bulb was blown out on Feb 4, 2017 after 2 months of operation.
At the sites, you probably know that we blow our spectrum out of the water with the calibration lines, with SNRs of about 100 on the scale of about 10 seconds. For us this might be impossible, since we aren't as quiet.
If we want 1% calibration on our sweeps, we'll need 0.01 = Uncertainty = sqrt( (1 - COH^2)/(2 * Navg * COH^2) ), where COH is the coherence of the transfer function measurement and Navg is the number of measurements at a specific frequency. This equation comes from Bendat and Piersol, and is subject to a bunch of assumptions which may not be true for us (particularly, that the plant is stationary in time).
If we let Navg = 10, then COH ~ 0.999.
Coherence = Gxy^2/(Gxx * Gyy), where x(t) and y(t) are the input signal and output signal of the transfer function measurement, Gxx and Gyy are the spectral densities of x and y, and Gxy is the cross-spectral density.
Usually SNR = P_signal / P_noise, but for us SNR = A_signal / A_noise.
Eric Q and Evan H helped me find the relationship between Coherence and SNR:
P = Pn + Pc, Pn = P * (1 - Coh), Pc = P * Coh
==> SNR = sqrt( Pc / Pn ) = sqrt( Coh / 1 - Coh )
From Coh ~ 0.999, SNR ~ 30.
Question for Craig: What does the SNR of our lines have to be? IF we're only trying to calibrate the actuator in the audio band over long time scales, it seems we could get by with more frequency noise. Assuming we want a 1% calibration at 50-500 Hz, what is the requirement on the frequency noise PSD curve?
Our new janitor Francisco is started working in IFO room today.
Large film crews are working just out side the north west corner of the lab. They started around ~ 5:30am Do not plan on working late tonight.
ETMX sus damping restored.
C1:PSL-FSS_RMTEMP, C1:PSL-PMC_PMCTRANSPD and C1:PEM-count_temp channels are not reading since Friday
Here is a comparison of the error signal spectra after increasing the IMC modulation depth, to the contribution with RF inputs / whitening inputs terminated (which I borrowed from Koji's characterization of the same in Dec 2016, these shouldn't have changed).
Some general observations:
I will update with the in-loop error signal spectra, which should give us some idea of the loop bandwidth.
I will look into lowering the sampling rate, and how much out-of-band power is aliasing into the 0-256 Hz band and update with my findings.
Yikes. Please change the all teh WFS DQ channels sample rates from 2048 down to 512 Hz. I doubt we ever need anything about 180 Hz.
There is sometimes an issue with this: if our digital AA filters are not strong enough, the noise about above 256 Hz can alias into the 0-256 Hz band. We ought to check this quantitatively and make some elog statement about our AA filters. This issue is also seen in DTT when requesting a low frequency spectrum: DTT uses FIR filters which are sometimes not sharp enough to prevent this issue.
Turns out the "problem" with WFS2 and the apparent offset accumulation on the IMC Servo board is probably a slow machine problem.
Today, Koji and I looked at the situation a little more closely. This anomalous behaviour of the C1:IOO-MC_SUM channel picking up an offset seems correlated with light being incident on WFS2 head. Placing an ND filter in front of WFS 2 slowed down the rate of accumulation (though it was still present). But we also looked at the in-loop error signal on the IMC board (using the "Out 2" BNC on the front panel), and this didn't seem to show any offset accumulation. Anyways, the ability of the Autolocker doesn't seem to be affected by this change, so I am leaving the WFS servo turned on.
The new demod phases (old +45degrees) and gains (old gains *0.2) have been updated in the SDF table. It remains to see that the WFS loops don't drag the alignment over longer timescales. I will post a more detailed analysis here over the weekend...
Also, we thought it would be nice to have DQ channels for the WFS error signals for analysis of the servo (rather than wait for 30 mins to grab live fine resolution spectra of the error signals with the loop On/Off). So I have added 16 DQ channels [recorded at 2048 Hz] to the c1ioo model (for the I and Q demodulated signal from each quadrant for the 8 quadrants). The "DRATE" for the c1ioo model has increased from ~200 to 410. Comparing to the "DRATE" of c1lsc, which is around 3200, we think this isn't significantly stretching the DAQ abilities of the c1ioo model...
Koji, Gautam, Johannes
We quickly checked the situation of the projector in the control room.
- We found that the proejctor was indicating "lamp error".
==> Steve, could you remove the projector from the ceiling and check if it still does not work?
If it still does not work, send it back to the vender. It should be covered by the previous service.
- Zita seemed happy with the DVI output. We tried the dual display configration and VGA and DVI are active right now.
The DVI output (from RADEON something video card) is somewhat strange. We probably need to look into the video display situation.
Craig and I have been trying to put together a Simulink diagram of the proposed alternative calibration scheme. Each time I talk the idea over with someone, I convince myself it makes sense, but then I try and explain it to someone else and get more confused. Probably I am not even thinking about this in the right way. So I am putting what I have here for comments/suggestions.
What's the general idea?
Suppose the PSL is locked to the MC cavity, and the AUX laser is locked to the arm cavity (with sufficiently high BW). Then by driving a line in the arm cavity length, and beating the PSL and AUX lasers, we can determine how much we are modulating the arm cavity length in metres by reading out the beat frequency between the two lasers, provided the arm cavity length is precisely known.
So we need:
To be able to sense a 1kHz line being driven at 1e-16 m amplitude, I estimate we need a beat note stability of ~1mHz/rtHz at 1kHz.
Requirements and what we have currently:
On the hardware side of things, we need:
Koji and I briefly looked through the fiber inventory we have yesterday. We have some couplers (one mounted) and short (5m) patch fibers. But I think the fiber infrastructure we have in place currently is adequate - we have the AUX light brought to the PSL table, and there is a spare fiber running the other way if we want to bring the PSL IR to the end as well.
I need to also think about where we can stick the EOM in given physical constraints on the EX table and the beam diameter/aperture of EOM...
Following the discussion at the meeting today, I wanted to finish up the WFS tuning and then hand over the IFO to Johannes for his loss stuff. So I did the following:
At this point, I figured I would leave the WFS in this state and observe its behaviour overnight. But abruptly, the IMC behaviour changed dramatically. I saw first that the IMC had trouble re-acquiring lock. Moreover, the PC Drive seemed saturated at 10.0V, even when there was no error signal to the MC Servo board. Looking at the MEDM screen, I noticed that the "C1-IOO_MC_SUM_MON" channel had picked up a large (~3V) DC offset, even with In1 and In2 disabled. Moreover, this phenomenon seemed completely correlated with opening/closing the PSL shutter. Johannes and I did some debugging to make sure that this wasn't a sticky button/slider issue, by disconnecting all the cables from the front panel of the servo board - but the behaviour persisted, there seemed to be some integration of the above-mentioned channel as soon as I opened the PSL shutter.
Next, I blocked first the MC REFL PD, and then each of the WFS - turns out, if the light to WFS2 was blocked and the PSL shutter opened, there was no integrating behaviour. But still, locking the MC was impossible. So I suspected that something was wrong with the LO inputs to the WFS Demod Boards. Sure enough, when I disconnected and terminated those outputs of the RF distribution box, I was able to re-lock the MC fine.
I can't explain this bizzare behaviour - why should an internal monitor channel of the MC Servo board integrate anything when the only input to it is the backplane connector (all front panel inputs physically disconnected, In1 and In2 MEDM switches off)? Also, I am not sure how my work on the WFS could have affected any hardware - I did not mess around at the 1X1 rack in the evening, and the light has been incident on the WFS heads for the past few days. The change in modulation depth shouldn't have resulted in the RF power in this chain crossing any sort of damage threshold since the measured power before the changes was at the level of -70dBm, and so should be at most -40dBm now (at the WFS demod board input). The only thing different today was that the digital inputs of the WFS servos were turned on...
So for tonight I am leaving the two outputs of the RF distribution box that serve as the LO for the WFS demod boards terminated, and have also blocked the light to both WFS with beam blocks. The IMC seems to be holding lock steady, PC drive levels look normal...
Unrelated to this work, but I have committed to the svn the updated versions of the mcup and mcdown scripts, to reflect the new gains for the autolocker...
This is already how it's hooked up. The hole on the from that says +24 V is for an indicator light.
The amplifier unit should use the three pin dsub connectors (3w3?) that we use on many of the other units for DC power, and preferably go through the back panel. You can leave out the negative pin, since you just need +24 and ground.
There has been a change in the default format for the output of the condor_q command at CIT clusters. This could be problematic for the summary page status monitor, so I have disabled the default behavior in favor of the old one. Specifically, I ran the following commands from the 40m shared account: mkdir -p ~/.condor echo "CONDOR_Q_DASH_BATCH_IS_DEFAULT=False" >> ~/.condor/user_config This should have no effect on the pages themselves.
I finished designing the PCBs for the VME crate back sides (see attached). The project files live on the DCC now at https://dcc.ligo.org/LIGO-D1700058. I ordered a prototype quantity (9) of the PCB printed and bought the corresponding connectors, all will arrive within the next two weeks. See also attached the front panels for the Acromag DAQ chassis and Lydia's RF amplifier unit (the lone +24V slot confuses me: I don't see a ground connector?). On the Acromag panel, six (3x2) of the DB37 connectors are reserved for VME hardware, two are reserve, and I filled the remaining space with general purpose BNC connectors for whatever comes up.
I made a tentative front panel design for the newly installed amplifier box. I used this chassis diagram to place the holes for attaching it. I just made the dimensions match the front of the chassis rather than extending out to the sides since the front panel doesn't need to screw into the rack; the chassis is mounted already with separate brackets. For the connector holes I used a caliper to measure the feedthroughs I'm planning to use and added ~.2 mm to every dimension for clearance, because the front panel designer didn't have their dimensions built in. Please let me know if I should do something else.
The input and coupled output will be SMA connectors since they are only going to the units directly above and below this one. The main output to the EOM is the larger connector with better shielded cables. I also included a hole for a power indicator LED.
EDIT: I added countersinks for 4-40 screws on all the screw clearance holes.
Johannes, if you're going to be putting a front panel order in soon, please include this one.
Also, Steve, I found a caliper in the drawer with a dead battery and the screws to access it were in bad shape- can this be fixed?
Brief Summary: I am currently looking at the acoustic noise around both arms to see if there are any frequencies from machinery around the lab that stand out and to see what we can remove/change. I am using a Bluebird microphone suspended with surgical tubing from the cable trays to isolate it from vibrations. I am also using a preamp and the SR875 spectrum analyzer taking 6 sets of data every 1.5 meters (0 to 200Hz, 200Hz to 400Hz, 400z to 800Hz, 800Hz to 3200Hz, 3.2kHz to 12kHz, 12kHz to 100kHz).
· Attachment 1 is a PSD of the first 3 measurements (from 0 to 12kHz) that I took every 1.5 meters along the x arm with the preamp and spectrum analyzer
· Attachment 2 is a blrms color map of the first 6 sets of data I took (from 2.4m to 9.9m)
· Attachmetn 3 is a picture of the microphone set up with the surgical tubing
Problems that occurred: settings on the preamp made the first set of data I took significantly smaller than the data I took with the 0dB button off and the last problem I had was the spectrum analyzer reading only from -50 to -50 dBVpk
Gautam and I were able to get the Raspberry Pi up and running today, including being able to ssh into it from the control room.
Below are some details about the setup/procedure that might be helpful to anyone trying to establish Ethernet connection for a new RPi, or a new operating system/SD card.
Here is the physical setup:
The changes that need to be made for a new Raspbian OS in order to communicate with it over ssh are as follows, with links to tutorials on how to do them:
1. Edit dhcpcd.conf file: https://www.modmypi.com/blog/how-to-give-your-raspberry-pi-a-static-ip-address-update
2. Edit interfaces file: https://www.mathworks.com/help/supportpkg/raspberrypi/ug/getting-the-raspberry_pi-ip-address.html
3. Enable ssh server: http://www.instructables.com/id/Use-ssh-to-talk-with-your-Raspberry-Pi/
The specific addresses for the RPi we set up today are:
IP Address: 192.168.113.107
Gateway/Routers/Domain Name Servers: 192.168.113.2
GV: I looked through /etc/var/bind/martian.hosts on chiara and decided to recycle the IP address for Domenica.martian as no RPis are plugged in right now... I'm also removing some of the attachments that seem to have been uploaded multiple times.
I was a little confused why the In1 Gain had to be as high as +10dB - before the changes to the RF chain, we were using +27dB, and we expect the changes made to have increased the modulation depth by a factor of ~25, so I would have expected the new In1 Gain to be more like 0dB.
While walking by the PSL table, I chanced upon the scope monitoring PMC transmission, and I noticed that the RIN was unusually high (see the scope screenshot below). We don't have the projector on the wall anymore, but it doesn't look like this has shown up in the SLOW monitor channel anyways. Disabling the MC autolocker / closing the PSL shutter had no effect. I walked over to the amplifier setup in 1X2, and noticed that the SMA cable connecting the output of the amplifier to the EOM drive was flaky. By touching the cable a little, I noticed that the trace on the scope appeared normal again. Turning off the 29.5MHz modulation source completely returned the trace to normal.
So I just made a new cable of similar length (with the double heat shrink prescription). The PMC transmission looks normal on the scope now. I also re-aligned the PMC for good measure. So presumably, we were not driving the EOM with the full +27dBm of available power. Now, the In1 Gain on the MC servo board is set to +2dB, and I changed the nominal FSS FAST gain to +18dB. The IMC OLTF now has a UGF of ~165kHz, though the phase margin is only ~27 degrees..
MC Servo Board
I would think that we want to fix the I/Q orthog inside the demod board by trimming the splitter. Mixing the Q phase signal to the I would otherwise allow coupling of low frequency Q phase junk from HOMs into the MC lock point.
Of course this doesn't matter for the IMC locking as we only use the I phase signal, but
29.5 MHz RF Modulation Source
IMC Demodulation Board
I wanted to do a quick check to see if the observed signal levels were in agreement with tests done on the workbench with the Marconi. The mixers used, JMS-1H, have an advertised conversion loss of ~7dB (may be a little higher if we are not driving the LO at +17dBm). The Lissajous ellipse above is consistent with these values. I didn't measure powers with the MC REFL PD plugged into the demod board, but the time series plot above suggest that I should have ~0dBm power in the MC REFL PD signal at 29.5MHz for the strongest flashes (~0.3Vpp IF signal for the strong flashes).
MC Servo Board
Some general remarks
The input impedance of the mixer is not constant. As the diode switches, it changes dynamically. Because of this, the waveform of the LO at the mixer input (i.e. the amplifier output) is not sinusoidal. Some of the power goes away to harmonic frequencies. Also, your active probe is calibrated to measure the power across the exact 50Ohm load, which is not in this case. The real confirmation can be done by swapping the mixer with a 50Ohm resistor. But it is too much. Just confirm the power BEFORE the amp is fine. +/-1dB does not change the mixer function much.
Instead, we should measure
- Gain imbalance
of the I/Q output. This can be checked by supplying an RF signal that is 100~1kHz away from the LO frequency and observe I&Q outputs.
Rana and I spent some time looking at the IMC demod board earlier today. I will post the details shortly, but there was a label on the front panel which said that the nominal LO level to the input should be -8dBm. The new 29.5MHz routing scheme meant that the LO board was actually being driven at 0dBm (that too when the input to the RF distribution box was attenuated by 5dB).
An elog search revealed this thread, where Koji made some changes to the demod board input attenuators. Rana commented that it isn't a good idea to have the LO input be below 0dBm, so after consulting with Koji, we decided that we will
After implementing these changes, and testing the board with a Marconi on the workbench, I found that the measured power levels (measured with an active FET probe) behave as expected, up till the ERA-5SM immediately prior to the LO (U4 and U6 on the schematic). However, the power after this amplifier (i.e. the input to the on-circuit LO, Minicircuits JMS-1H, which we want to be +17dBm), is only +16dBm. The input to these ERA-5SMs, which are only ~2years old, is -2dBm, so with the typical gain of +20dB, I should have 18dBm at their output. Moreover, increasing the input power to the board from the Marconi doesn't linearly increase the output from the ERA-5SM. Just in case, I replaced one of the ERA-5SMs, but observed the same behaviour, even though the amplifier shouldn't be near saturation (the power upstream of the ERA-5SM does scale linearly).
This needs to be investigated further, so I am leaving the demod board pulled out for now...
To remind myself about how to put filter caps on the mini-circuits RF Amps, I looked at Koji's recent elog. Its mostly about op-amps, but the idea holds for us.
We want a big (~100 uF) electrolytic with a 50V rating for the +24V RF Amp. And then a 50V ceramic capacitor of ~0.1 uF close to the pins. Remember that the power feed through on the Mini-circuits case is itsself a capacitive feedthrough (although I guess its a ~100 pF).
Later, we should install in this box an active EMI filter (e.g. Vicor)
c1iool0 was down again. Rather than key the crate, this time I just pushed the reset button on the front and it came back.
As move towards the wonderfulness of AcroMag, we also have to buy a computer to handle all of these IOCs. Let's install the new c1iool0 over by the SUS computer.
To install the replacement amplifier, I did the following:
Still to be done:
Lydia finished up installing the new RF amplifier, and will elog the details of the installation.
I wanted to try and measure the IMC OLG to compare against my Simulink model. So I went about performing a few checks. Summary of my findings:
TBC tomorrow, I'm leaving the PSL shutter closed and the RF source off for tonight...
Here are the details as promised.
Attachment #1: Updated simulink model. Since I haven't actually run this model, all the TF blocks are annotated "???", but I will post an updated version once I have run the model (and fix some of the questionable aesthetic choices)
Attachment #2: Measured and fitted transfer functions from the "IN1" input (where the demodulated MC REFL goes) to the "SERVO" output of the MC servo board (to FSS box). As mentioned in my previous elog, I had to put in a pole (fitted to be at ~2MHz, called pole 9 in the plot) in order to get good agreement between fit an measurement up to 10MHz. I didn't bother fitting all the high frequency features. Both gain sliders on the MEDM screen ("IN1 Gain" and "VCO gain") were set to 0dB for this measurement, while the super boosts were all OFF.
Attachment #3: Measured and fitted transfer function from "TEST 1 IN" to "FAST OUT" of the FSS box. Both gains on the FSS MEDM screen ("Common gain adjust" and "fast gain adjust") were set to 0dB for this measurement. I didn't need any ad-hoc poles and zeros for this fit (i.e. I can map all the fitted poles and zeros to the schematic), but the fit starts to deviate from the measurement just below 1 MHz.. perhaps I need to add a zero above 1MHz, but I can't see why from the schematic...
Attachment #4: Measured TF from "TEST 1 IN" to "PC OUT" on the FSS box. MEDM gains were once again 0dB. I can't get a good fit to this, mainly because I can't decipher the poles and zeros for this path from the schematic (there are actually deviations from the schematic posted on the 40m DCC page in terms of component values, I will try and correct whatever I notice) . I'll work on this...
Attachment #5: Data files + .fil files used to fit the data with LISO
Data + plots + fits + updated schematics to follow...
Most of the model has come together, I am not too far from matching the modelled OLG to the measured OLG. So I will now start thinking about designing the controller for the MCL part (there are a couple of TFs that have to be measured for this path).
I'd like to fix a few things at 1X1 when we plug in the new amplifier for the 29.5MHz modulation signal.
Steve has ordered rolls of pre-twisted wire to run from 1X1 to the PSL table, so that part can be handled later.
But at 1X1, we need to tap new paths from +/- 24V to the DIN connectors. I think it's probably fine to turn off the two Sorensens, do the wiring, and then turn them back on, but is there any procedure for how this should be done?
Quick summary elog, details to follow. I did the following:
The measurements I have look reasonable. But I had a hard time trying to look at the schematic and determine what is the appropriate number and locations of poles/zeros with which to fit the measured transfer function. Koji and I spent some time trying to go through the MC Servo board schematic, but looks like the version uploaded on the 40m DCC tree doesn't have changes made to it reflected (we compared to pictures on the 40m google photos page and saw a number of component values were different). Since the deviation between fit and measurement only occurs above 1MHz (while using poles/zeros inferred from the schematic), we decided against pulling out the servo board and investigating further - but this should be done at the next opportunity. I've marked the changes we caught on a schematic and will upload it to the 40m DCC page, and we can update this when we get the chance.
So it remains to fit the other two measured TFs, and add them to the Simulink model. Then the only unknown will be the PDH discriminant, which we anyway want to characterize given that we will soon have much more modulation.
The ITMX oplev beam is clipping. It will be corected with locked arm
After fighting with Altium for what seems like an eternity I have finished putting my vision of the vme crate backplane adapter board into an electronic format. It is dimensioned to fill the back space of the crate exactly. The connectors are panel mount and the PCB attaches to the connectors with screws, such that the whole thing will be mechanically much more stable than the current configuration. A mounting bracket will attach to horizontal struts that need to be installed in the crates, mechanical drawings to follow.
I set everything up and connected it as shown on the block diagram attached to the previous entry, with the exception of the DC power. This is becuase there is no place open to connect to on the DIN rail where the DC power is distributed, so the +24V power will have to be shut off to the other equipment in 1X1 before we can connect the amplifier. (The amplifier is in 1X2, but the DC power distribution was more accessible in 1X1.) I also added 3 new +24 V clips with fuses despite needing only one, so next time we need to connect something new it's not such a hassle.
The RF distribution box where the 29.5 MHz signal originates should not be turned on until the amplifer has DC power. Since we may have a power interruption tomorrow, the plan is to wait until things are shut down in preparation, and then shut off anyhting else necessary before connecting the new clips on the rail to the existing ones.
received this note: at 4:11pm Tuesday, Feb 7, 2017
Date: Wednesday, February 8, 2017
Time: 7:30 AM – 8:30 AM
Contact: Rick Rodriguez x-2576
Pasadena Water and Power (PWP) will be performing a switching operation of the
Caltech Electrical Distribution System that is expected to be transparent to Caltech,
but could result in a minor power anomaly that might affect very sensitive equipment.
IMPACT: Negligible impact......?
There may be temporary power interruption tomorrow!
PS:we did not see any effect